OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=2, b=5, c=?

OpenStudy (anonymous):

do you know the Pythagorean theorem?

OpenStudy (anonymous):

a\[a ^{2}+b^2=c^2\]

OpenStudy (anonymous):

yess i do

OpenStudy (anonymous):

so can you solve it with that equation?

OpenStudy (anonymous):

\[2^2+5^2=c^2\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

thank you